### v1

# A study of bound states of $\eta,\eta^{{}^{\prime}},D^{0}$, $\bar{D^{0}}$, $B^{0}$, $\bar{B^{0}}$, $\bar{K^{0}}$ and $\phi$ mesons with light and heavy nuclei within chiral SU(3) model

Abstract
In the present work we explore the possibilities of the formation of bound states of neutral pseudoscalar mesons $\eta,\eta^{{}^{\prime}},D^{0}$, $\bar{D^{0}},B^{0},\bar{B^{0}}$ and $\bar{K^{0}}$ and the vector meson $\phi$, with the nuclei 12C, 16O, 40Ca and 208Pb, calculating their binding energy and absorption decay width. To calculate the optical potentials of these mesons in different nuclei under study, we shall use the chiral SU(3) hadronic mean field model, in which the properties of nucleons in the medium are modified through the scalar isoscalar fields $\sigma$ and $\zeta$ and the scalar-isovector field $\delta$. The scalar-isovector field $\delta$ account for the finite isospin asymmetry of different nuclei having asymmetry in number of protons and neutrons. The binding energy and absorption decay width of mesons are calculated for the ground state and some of possible excited states of the nuclei. In the chiral SU(3) model, the mesons $\eta,D^{0}$, $\bar{D^{0}},B^{0},\bar{B^{0}}$ and $\bar{K^{0}}$ are observed to have significant negative mass shift upto nuclear saturation density which lead to the possibility of the bound states at least for ground states and some of excited states in case of heavy nuclei. For the pseudoscalar singlet $\eta^{{}^{\prime}}$ and the vector meson $\phi$ the mass shift obtained are found to be small and bound states are not formed. The present calculations are compared with different studies available in the field and will be useful in understanding the outcomes from different experimental facilities focusing on this area of research.
I Introduction
The theory of strong interactions, quantum chromodynamics (QCD), has two well known properties: asymptotic freedom and confinement. The understanding of the low-energy non-perturbative regime of QCD poses extra challenges due to large value of coupling constant. The properties of QCD such as breaking of chiral symmetry and its expected restoration at high temperature and baryon density play an essential role in our understanding of strong interaction physics. The study of medium modification of hadron properties, specially pseudoscalar and vector mesons, play a pivotal role in revealing the nature of the partial restoration of the chiral symmetry at high temperature and baryon densities. The heavy-ion collision experiments, for example, the large hadron collider (LHC) at CERN, Switzerland and relativistic heavy-ion collider (RHIC) at BNL, USA, explore the regime of strongly interacting matter at high temperature and low baryon densities, whereas, the future experimental facilities such compressed baryon matter (CBM) of FAIR, GSI and the Nuclotron-based Ion Collider fAcility (NICA) at Joint Institute for Nuclear Research (JINR), Russia intend to explore the nature of interactions and phase transitions at high baryon densities and low/moderate temperatures. There are extensive theoretical studies using various approaches which explore the medium modification of hadron properties at finite temperature and baryon densities [1, 3, 2, 4].
An alternative technique which play a significant role in understanding the strongly interacting matter and role of chiral symmetry is the study of mesic nuclei, i.e., the bound states of mesons and nuclei, where strong interactions play a dominant role. In the formation of mesic nuclei, one of the nucleon in a given nuclei is replaced with the meson. This is complementary to mesic atoms in which mesons replaces electron in the atom and Coulomb interactions play the key role [5]. Formation of mesic nuclei for charged and neutral mesons has been studied using different approach, for example, quark meson coupling (QMC) model [7, 6], Nambu–Jona-Lasinio model (NJL) [8], chiral unitary approach [9, 10] etc. The mesons which have strong attractive interactions in the dense nuclear medium, i.e., whose mass deceases with increase in the baryon density and thus, have deep attractive potential, may form the bound states with different nuclei.
The study of $\eta$ and $\eta^{{}^{\prime}}$ mesic nuclei shed light not only on interactions of these mesons with nucleons in the nuclear matter, but is also important to deepen our understanding on the behaviour of $U_{A}(1)$ anomaly in the dense medium. Initial studies in the direction of $\eta$ mesic nuclei were initiated by Haider and Liu [11]. In this work, through the analysis of $\pi N\longrightarrow\eta N$ cross-sections and solving relativistic wave equation, possibilities of existence of bound states of $\eta$ mesons with nuclei having mass number $A\geq 12$ have been claimed. In [12], Green function method is applied to study the $\eta-^{16}$ bound state. Subsequent to these initial studies, in-medium properties of $\eta$ and $\eta^{{}^{\prime}}$ mesons and possibility of their bound state formation with different nuclei have been studied by different groups [13, 19, 14, 16, 17, 18]. The formation of $\eta^{{}^{\prime}}(958)-$nucleus bound state is investigated in the $(p,d)$ reaction to explore the medium modification of $\eta^{{}^{\prime}}(958)$ mass [13]. Many experiments are devoted to search the $\eta$ mesic nuclei by different collaborations by studying reactions from heavy and light targets. Experiments at BNL [20] involved the reaction of $\pi^{+}$ with targets lithium, carbon, oxygen, and aluminium, whereas, Los Alamos Meson Physics Facility (LAMPF) considered 18O [21], but both with negative results. The photo-production process, induced by reaction of $\gamma$ photon on 12C, lead to the formation of $\eta-$nucleon bound state, as claimed by Lebedev Physical Institute (LPI) group [22]. Deuteron induced reaction on carbon nuclei was studied in JINR to investigate the formation of $\eta$ mesic nuclei through the decay of $N^{*}(1535)$ resonance [23]. Along with heavy nuclei as targets, experimentalist are actively looking for the bound state of $\eta-$ mesons with light nuclei, for example, WASA at COSY collaboration through reaction channels $pd\rightarrow^{3}$e 2 $\gamma$ and $pd\rightarrow^{3}$e 6 $\gamma$ are looking for bound states of $\eta$ with 3He nuclei [24, 25]. Theoretically, using few body $\eta-NNN$ system existence of 3He $\eta$ bound state is explored [26]. Recent reviews on the theoretical and experimental efforts on $\eta$ and $\eta^{{}^{\prime}}$ mesic nuclei are presented in Refs. [28, 27].
Investigation of the medium modification of open charm $D$ mesons have implications in understanding the their prouction rate, the phenomenon of charmonium suppression which may shed light on formation of QGP, and also in exploring the possibility of formation of $D$ [10, 29] and $J/\psi$ mesic nuclei [30]. As the open charm $D$ mesons have one light $u/d$ quark or anti-quark as its content, these mesons can undergo the significant mass modifications, through the medium modification of light quark condensates. The medium modifications of charm mesons have been studied using QCD sum rules, chiral SU(3) hadronic model, coupled channel approach, QMC model etc. [31, 32, 33, 34, 38, 35, 36, 37, 39]. In Ref. [7], using the QMC model, bound states of $D^{-},D^{0}$ and $\bar{D^{0}}$ mesons in 208Pb were studied. Using the $D$ meson self-energies, calculated using the coupled channel approach, the bound states of $D^{0}$ mesons with 12C and 208Pb were studied in Ref. [10]. The study of $D$ meson bound states with the nuclei is also importrant for the $\bar{\text{P}}$ experiment of FAIR project [62]. In literature the work has also been done on the bound states of $\omega$ [40], $J/\psi$ [30], $\eta_{c}$ [41] and $\eta_{b}$ [42] mesons with different nuclei. Although the in-medium mass of bottom meson have been studied using different approaches [43], the $B-$mesic nuclei have not been explored except one recent study using QMC model [44].
Understanding the medium modification of $\phi$ mesons is also an active topic at present due to different results obtained from various experimental studies. In the KEK-PS E325 experiment, though the study of invariant mass spectra of $e^{-}e^{+}$ pairs produced in 12 GeV $p+A$ reaction, through the process $p+A\rightarrow\phi+X\rightarrow e^{-}e^{+}+X^{{}^{\prime}}$, a negative mass shift of $3.4$% and decay width $14.5$ MeV is observed at nuclear saturation density [45]. Other experiments, for example, Laser-Electron Photon facility at SPring-8 (LEPS) [46] and by CLAS collaboration JLab [47, 48], have reported significant modifications in the in-medium decay width, without significant effect on mass shift. The J-PARC E16 collaboration aim to investigate the mass modifications of $\phi$ mesons through the mass spectra of $e^{-}e^{+}$ pairs, as in KEK, but with more high statistics and in the energy range of 50 GeV [49, 50]. Theoretically, the medium modifications of $\phi$ mesons have been studied using QCD sum rules, coupled channel approach, chiral hadronic model, quark meson model etc. [51, 52, 54, 53, 55, 6, 56, 57, 58, 59]. The negative mass shift for the $\phi$ mesons observed under various theoretical approaches and also as reported by KEK-PS 325 [45], motivate to study the possibility of the formation of $\phi-$ mesic nuclei [6]. At JPARC experimental facility, plans are to investigate the mass shift of $\phi$ mesons though the $\phi$ meson-nuclei bound states using reactions $\bar{p}p\rightarrow\phi\phi$ [60, 61].
In the present paper, we shall explore the formation of bound states of pseudoscalar mesons, $\eta,\eta^{{}^{\prime}},D^{0},\bar{D^{0}},B^{0},\bar{B^{0}},\bar{K^{0}}$ and the vector meson $\phi$, whereas, the bound states of $\omega$, $J/\psi$, $\eta_{c}$ and $\eta_{b}$ will be investigated in future work. For our current work, we shall use the chiral SU(3) hadronic mean field model which is based on low energy properties of QCD [63]. Chiral SU(3) model has been used in literature to study the modification of properties of kaons [65, 66, 64], vector mesons ($\omega,\rho$ and $\phi$) [67], pseudoscalar $\eta$ meson [68, 69], $D$ [38, 70, 39] and $B$ mesons [71], charmonium [72] and bottomonium [73], in nuclear and strange hadronic matter. The optical potentials of mesons under study, calculated using chiral SU(3) for different light and heavy nuclei, will be used to solve Klein Gordon equation in the momentum space to obtain the binding energies and decay width.
Following is the outline of present paper: In Sec. II, we shall present the details of chiral SU(3) model. Details of how to obtain the optical potentials of mesons in the environment of nuclei will be discussed in Sec. III. The results and discussion will be presented in Sec. IV and finally, in Sec. V, summary and conclusion will be given.
II Chiral SU(3) hadronic model
To investigate the formation of mesic nuclei in the present work, we shall use the chiral SU(3) hadronic mean-field model based on nonlinear realization of chiral symmetry [63, 74, 75, 76]. In the chiral SU(3) model, the interactions between the nucleons are mediated through the exchange of non-strange scalar meson $\sigma$, strange scalar meson $\zeta$, scalar-isovector meson $\delta$, vector meson $\omega$ and vector-isovector meson $\rho$. The scalar field $\delta$ and vector field $\rho$ contribute when medium has finite isospin asymmetry. In this model, the scalar dilaton field, $\chi$, known as glueball field is introduced to incorporate the broken scale invariance property of the QCD [63]. The Lagrangian density for chiral SU(3) hadronic mean-field model is written as [63]
$\mathcal{L}_{\text{chiral}}$ $=\mathcal{L}_{\text{kin}}+\sum_{M=P,X,V,A}\mathcal{L}_{\text{BM}}+\mathcal{L}_{0}+\mathcal{L}_{\text{vec}}+\mathcal{L}_{\text{SB}}.$ (1)
In Eq. (1), $\mathcal{L}_{\text{kin}}$ represents the kinetic energy term for baryons and mesons and is given by
$\mathcal{L}_{\text{kin }}$ $=i\operatorname{Tr}\bar{B}\gamma_{\mu}D^{\mu}B+\frac{1}{2}\operatorname{Tr}D_{\mu}XD^{\mu}X+\operatorname{Tr}\left(u_{\mu}Xu^{\mu}X+Xu_{\mu}u^{\mu}X\right)\quad+\frac{1}{2}\operatorname{Tr}D_{\mu}YD^{\mu}Y$
$+$ $\frac{1}{2}D_{\mu}\chi D^{\mu}\chi-\frac{1}{4}\operatorname{Tr}\left(V_{\mu\nu}V^{\mu\nu}\right)-\frac{1}{4}\operatorname{Tr}\left(\mathcal{A}_{\mu\nu}\mathcal{A}^{\mu\nu}\right).$ (2)
In Eq. (2), the first term represents the kinetic energy term for the baryon octet, $B$. The covariant derivative $D_{\mu}$ appearing in this term is defined as $D_{\mu}B=\partial_{\mu}B+i\left[\Gamma_{\mu},B\right]$, with $\Gamma_{\mu}=-\frac{i}{4}\left[u^{\dagger}\partial_{\mu}u-\partial_{\mu}u^{\dagger}u+\right.$ $\left.u\partial_{\mu}u^{\dagger}-\partial_{\mu}uu^{\dagger}\right]$. Here, $u=\exp\left[\frac{i}{\sigma_{0}}\pi^{a}\lambda^{a}\gamma_{5}\right]$ is the unitary transformation operator through which pseudoscalar mesons enter into the calculations. The second and third terms of Eq. (2) are the kinetic terms for the scalar and pseudoscalar mesons. Fourth and fifth terms represent the kinetic terms for the pseudoscalar singlet $Y$ and the dilaton field $\chi$, respectively. The last two terms defined in terms of field tensors, $V_{\mu\nu}$ and $\mathcal{A}_{\mu\nu}$, represent the kinetic term of spin-1 vector and axial vector mesons, respectively.
Within the chiral SU(3) model, to describe the interactions of nucleons through the exchange of scalar fields $\sigma,\zeta$ and $\delta$ and vector fields $\omega$ and $\rho$, under the mean-field approximation, following interaction Lagrangian is written
$\mathcal{L}_{BM}=-\sum_{i}\overline{\psi_{i}}\left[g_{\omega i}\gamma_{0}\omega+g_{\rho i}\gamma_{0}\tau_{3}\rho+m_{i}^{*}\right]\psi_{i},$ (3)
where $i=p,n$. In the above, $m_{i}^{*}$ is the effective mass of nucleons defined as
$m_{i}^{*}=-\left(g_{\sigma i}\sigma+g_{\zeta i}\zeta+g_{\delta i}\tau_{3}\delta\right).$ (4)
The coupling constants $g_{\sigma i}$, $g_{\zeta i}$ and $g_{\delta i}$ describe the strength of interactions of nucleons with scalar fields. The term $\mathcal{L}_{\text{0}}$ of Eq. (1) gives the self interaction of scalar mesons $\sigma,\zeta$ and $\delta$ and also the interactions for dilaton field $\chi$. This term is written as
$\mathcal{L}_{\text{0}}$ $=-\frac{1}{2}k_{0}\chi^{2}(\sigma^{2}+\zeta^{2}+\delta^{2})+k_{1}(\sigma^{2}+\zeta^{2}+\delta^{2})^{2}+k_{2}(\frac{\sigma^{4}}{2}+\frac{\delta^{4}}{2}+\zeta^{4}+3\sigma^{2}\delta^{2})$
$+k_{3}\chi(\sigma^{2}-\delta^{2})\zeta-k_{4}\chi^{4}-\frac{1}{4}\chi^{4}\ln\frac{\chi^{4}}{\chi_{0}^{4}}+\frac{\delta}{3}\chi^{4}\ln\frac{\left(\sigma^{2}-\delta^{2}\right)\zeta}{\sigma_{0}^{2}\zeta_{0}}.$ (5)
In Eq. (5), last two terms account for scale breaking effects introduced in the chiral SU(3) model through the dilation field $\chi$. In Eq. (1), the term $\mathcal{L}_{\text{vec }}$ describes the self interactions of vector mesons through the Lagrangian density
$\mathcal{L}_{\text{vec }}$ $=\frac{1}{2}\frac{\chi^{2}}{\chi_{0}^{2}}\Big{(}m_{\omega}^{2}\omega^{2}+m_{\rho}^{2}\rho^{2}\Big{)}+g_{4}(\omega^{4}+6\omega^{2}\rho^{2}+\rho^{4}).$ (6)
The term, $\mathcal{L}_{SB}$, of Eq. (1) is the explicit symmetry breaking term and is given by
$\mathcal{L}_{SB}$ $=-\left(\frac{\chi}{\chi_{0}}\right)^{2}\left[m_{\pi}^{2}f_{\pi}\sigma+\left(\sqrt{2}m_{K}^{2}f_{K}-\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\right)\zeta\right].$ (7)
At zero temperature, the thermodynamic potential per unit volume of the grand canonical ensemble is expressed as
$\begin{array}[]{r}\frac{\Omega}{V}=-\sum_{i}\frac{\gamma_{i}}{(2\pi)^{3}}\int_{0}^{k_{F,i}}d^{3}\mathbf{k}\left[E_{i}^{*}(\mathbf{k})-\mu_{i}^{*}\right]-\mathcal{L}_{0}-\mathcal{L}_{\text{vec }}-\mathcal{L}_{\mathrm{SB}}-\mathcal{V}_{vac},\end{array}$ (8)
where $E_{i}^{*}(\mathbf{k})=\sqrt{\mathbf{k}^{2}+m_{i}^{*2}}$ and $\mu_{i}^{*}=\mu_{i}-g_{\omega i}\omega-g_{\rho i}\tau_{3}\rho$ In addition, the vacuum potential energy, $\mathcal{V}_{\text{vac}}$, is subtracted from Eq. (8) to achieve vanishing vacuum energy. The equations of motion for the non-strange $\sigma$, the strange scalar $\zeta$, the scalar isovector $\delta$, the vector $\omega$, the vector-isovector $\rho$, and the scalar dilaton $\chi$ fields are derived by minimizing the thermodynamic potential and are written as
$k_{0}\chi^{2}\sigma-2k_{2}\left(\sigma^{3}+3\sigma\delta^{2}\right)-2k_{3}\sigma\zeta\chi-4k_{1}\left(\sigma^{2}+\delta^{2}+\zeta^{2}\right)\sigma$
$-\frac{d}{3}\chi^{4}\left(\frac{2\sigma}{\sigma^{2}-\delta^{2}}\right)+\left(\frac{\chi}{\chi_{0}}\right)^{2}m_{\pi}^{2}f_{\pi}-\sum_{i}g_{\sigma i}\rho_{i}^{s}=0,$ (9)
$k_{0}\chi^{2}\zeta-4k_{2}\zeta^{3}+k_{3}\chi\left(\delta^{2}-\sigma^{2}\right)-4k_{1}\left(\sigma^{2}+\delta^{2}+\zeta^{2}\right)\zeta$
$-\frac{d}{3}\frac{\chi^{4}}{\zeta}+\left(\frac{\chi}{\chi_{0}}\right)^{2}\left[\sqrt{2}m_{K}^{2}f_{K}-\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\right]-\sum_{i}g_{\zeta i}\rho_{i}^{s}=0,$ (10)
$k_{0}\chi^{2}\delta+2k_{3}\chi\delta\zeta-2k_{2}\left(\delta^{3}+3\sigma^{2}\delta\right)-4k_{1}\left(\sigma^{2}+\delta^{2}+\zeta^{2}\right)\delta$
$+\frac{2}{3}d\chi^{4}\left(\frac{\delta}{\sigma^{2}-\delta^{2}}\right)-\sum_{i}g_{\delta i}\tau_{3}\rho_{i}^{s}=0,$ (11)
$\left(\frac{\chi}{\chi_{0}}\right)^{2}m_{\omega}^{2}\omega+g_{4}\left(12\rho^{2}\omega+4\omega^{3}\right)-\sum_{i}g_{\omega i}\rho_{i}=0,$ (12)
$\left(\frac{\chi}{\chi_{0}}\right)^{2}m_{\rho}^{2}\rho+g_{4}\left(12\omega^{2}\rho+4\rho^{3}\right)-\sum_{i}g_{\rho i}\tau_{3}\rho_{i}=0,$ (13)
and
$k_{0}\chi\left(\sigma^{2}+\delta^{2}+\zeta^{2}\right)+\left(4k_{4}-d\right)\chi^{3}+k_{3}\left(\delta^{2}-\sigma^{2}\right)\zeta+\chi^{3}\left[1+\ln\left(\frac{\chi^{4}}{\chi_{0}^{4}}\right)\right]$
$-\frac{4}{3}d\chi^{3}\ln\left(\left(\frac{\left(\sigma^{2}-\delta^{2}\right)\zeta}{\sigma_{0}^{2}\zeta_{0}}\right)\left(\frac{\chi}{\chi_{0}}\right)^{3}\right)+\frac{2\chi}{\chi_{0}^{2}}\left[m_{\pi}^{2}f_{\pi}\sigma+\left(\sqrt{2}m_{K}^{2}f_{K}-\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\right)\zeta\right]$
$-\frac{\chi}{\chi_{0}^{2}}\left(m_{\omega}^{2}\omega^{2}+m_{\rho}^{2}\rho^{2}\right)=0,$ (14)
respectively. In the present work, we shall discuss the formation of bound states for the nuclei 12C, 16O, 40Ca and 208Pb. As we shall describe later, the optical potentials of mesons required as input in the Klein Gordan equation for the study of bound states, will depend upon the scalar fields $\sigma,\zeta$ and $\delta$. For a nuclei with radius $R$, the values of scalar fields are required as a function of radial coordinator $r$, such that $r$ varies from zero to $R$. For this the coupled equations of motion for scalar and vector fields are solved for the baryon densities corresponding to these nuclei as a function of $r$. In the present work, we have considered the radial dependence of vector densities of nucleons through the parameterization in the form of harmonic oscillator or two parameter Fermi distribution function, as discussed below, and the second order derivative term of the form $-\nabla^{2}{\cal\sigma}(r)$ in the field equations given above is not considered.
Total baryon density, $\rho_{B}$, in a given nuclei is sum of densities of protons and neutrons, i.e.,
$\rho_{B}(r)=\sum_{i=p,n}\rho_{i}(r).$ (15)
For the nuclei upto ${}^{18}O$, we shall use the harmonic oscillator type density distribution, given as [77]
$\rho_{i}(r)=\rho_{i,0}\left(1+a_{i}\left(\frac{r}{R_{i}}\right)^{2}\right)\exp\left[-\left(\frac{r}{R_{i}}\right)^{2}\right].$ (16)
On the other side, for heavy nuclei, two parameter Fermi distribution function will be used in the present work [77] and is written as
$\rho_{i}(r)=\frac{\rho_{i,0}}{1+\exp\left[\frac{(r-R_{i})}{a_{i}}\right]}.$ (17)
In Eqs. (16) and (17), $\rho_{i,0}$ is the density of nucleons at the center of nuclei and $R_{i}$ and $a_{i}$ are the radii and diffuseness parameters, respectively corresponding to nucleons of type $i$. Values of these parameters for different nuclei are tabulated in Table I of Ref. [77]. For the nuclei with mass number $A$ and having different number of protons, $Z$ and neutrons, $N$, the isospin asymmetry parameter $I$ is given by, $I=(N-Z)/A$, which is further defined in terms of densities of nucleons as, $I=\frac{\rho_{n}(r)-\rho_{p}(r)}{2\rho_{B}(r)}$. The system of equations given by Eqs. (9) to (14) are solved for the densities in a given nuclei from center to surface, as given by Eq. (15). At zero temperature, the number density $\rho_{i}$ and the scalar density $\rho_{i}^{s}$ are related to the Fermi momentum $k_{F,i}$ through the relations
$\rho_{i}=\frac{\gamma_{i}k_{F,i}^{3}}{6\pi^{2}},$ (18)
and
$\rho_{i}^{s}=\gamma_{i}\int_{0}^{k_{F,i}}\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}\frac{m_{i}^{*}}{E_{i}^{*}(\mathbf{k})}=\frac{\gamma_{i}m_{i}^{*}}{4\pi^{2}}\left[k_{F,i}E_{i}^{*}-m_{i}^{*2}\ln\left(\frac{k_{F,i}+E_{i}^{*}}{m_{i}^{*}}\right)\right],$ (19)
respectively. In the above, $\gamma_{i}$ is the degeneracy factor for the nucleons.
III Optical potentials of mesons
To obtain the binding energy and absorption decay width of the mesic nuclei, we shall solve the Klein Gordon equation in the presence of local potential $V(r)$ and is written as
$\left(-\nabla^{2}+\left(\mu+V(r)\right)^{2}\right)\psi(r)=\epsilon^{2}\psi(r).$ (20)
In the above equation, $\mu$ is the reduced mass of meson and nuclei under study. The local potential $V(r)$ is complex in nature, i.e.,
$V(r)=U(r)-\frac{i}{2}W(r),$ (21)
where the real part $U(r)$ is related to the mass shift of mesons, whereas, the imaginary part $W(r)$ relates to the absorption of mesons in the nuclei under consideration. For the real part, $U(r)$, we write [5]
$U(r)=m_{\psi}^{*}(r)-m_{\psi}=\Delta m_{\psi}(\rho_{0})\frac{\rho_{B}(r)}{\rho_{0}},$ (22)
where $\rho_{0}$ denotes the baryon density at the center of nuclei and $\Delta m_{\psi}(\rho_{0})$ $[\psi=\eta,\eta^{{}^{\prime}},D^{0},\bar{D^{0}},B^{0},$ $\bar{B^{0}},\bar{K^{0}},\phi]$ is the mass shift at this density. Also, $\rho_{B}(r)$ is the baryon density distribution inside the nuclei and is given by Eq.(15). The potential $W(r)$ is related to the decay width $\Gamma_{0}$ through the relation [5]
$W(r)=\Gamma_{0}(\rho_{0})\frac{\rho_{B}(r)}{\rho_{0}}.$ (23)
In the present work, the value of $\Gamma_{0}(\rho_{0})$ for pseudoscalar mesons, $\eta,\eta^{{}^{\prime}},D^{0},\bar{D^{0}},B^{0},\bar{B^{0}}$ and $\bar{K^{0}}$ will be calculated following the procedure adopted in the calculations within QMC model [14]. Under this, $\Gamma_{0}(\rho_{0})$ is related to the mass shift $\Delta m_{\psi}(\rho_{0})$ through a parameter $\kappa$ using relation
$\Gamma_{0}(\rho_{0})=-\kappa\Delta m_{\psi}(\rho_{0})+\Gamma_{\text{vac}}.$ (24)
The parameter $\kappa$ imitate the absorption of mesons in the nuclear medium and $\Gamma_{\text{vac}}$ is the decay width in the vacuum. In our present work, the contributions from the first term of above equation will be considered to estimate the contribution of imaginary part of potential to the binding energy. We shall solve the Klein Gordon equation in the momentum space using the Fourier transformation and following the partial wave decomposition method as described in Refs. [78, 79] to obtain the complex energy eigen values $\epsilon$ which are further related to the binding energy $\cal E_{B}$ and decay width $\Gamma$ through relations $\cal E_{B}=\text{Re}\,\epsilon-\mu$ and $\Gamma=-2\,\text{Im}\,\epsilon$, respectively. In the following, we briefly describe the calculations of mass shift of $\eta,\eta^{{}^{\prime}},D^{0},\bar{D^{0}},B^{0},\bar{B^{0}}$, $\bar{K^{0}}$ and $\phi$ mesons which will be used as input to obtain the optical potentials of mesons in the environment of different nuclei.
III.1 Pseudoscalar $\eta$ and $\eta^{{}^{\prime}}$ mesons
In this section, we present the details of interaction Lagrangian density and dispersion relations to obtain the in-medium masses of pseudoscalar $\eta$ and $\eta^{\prime}$ mesons in the nuclear medium relevant for different nuclei. The physical states $\eta$ and $\eta^{\prime}$ result due to the mixing of $\eta_{8}$ and $\eta_{0}$ states. The pseudoscalar $\eta_{8}$ appears in the octet of pseudoscalar mesons defined by [63]
$P=\frac{1}{\sqrt{2}}\pi_{a}\lambda^{a}=\left(\begin{array}[]{ccc}\frac{1}{\sqrt{2}}\left(\pi^{0}+\frac{\eta_{8}}{\sqrt{1+2w^{2}}}\right)&\pi^{+}&2\frac{K^{+}}{w+1}\\ \pi^{-}&\frac{1}{\sqrt{2}}\left(-\pi^{0}+\frac{\eta_{8}}{\sqrt{1+2w^{2}}}\right)&2\frac{K^{0}}{w+1}\\ 2\frac{K^{-}}{w+1}&2\frac{K^{0}}{w+1}&-\frac{\eta_{8}\sqrt{2}}{\sqrt{1+2w^{2}}}\end{array}\right),$ (25)
where $w=\sqrt{2}\sigma_{0}/\zeta_{0}$. For the case of pseudoscalar single $\eta_{0}$, we define
$Y=\frac{1}{\sqrt{3}}\eta_{0}I.$ (26)
Within the chiral SU(3) model, for the pseudoscalar $\eta_{8}$, the interaction Lagrangian density is written as
$\mathcal{L}_{{\eta_{8}}B}$ $=$ $-\frac{1}{2}\left(m_{\eta_{8}}^{2}-\frac{(\sqrt{2}\sigma^{\prime}-4\zeta^{\prime})m^{2}_{\pi}f_{\pi}+8\zeta^{\prime}m^{2}_{K}f_{K}}{\sqrt{2}f^{2}}\right)\eta_{8}^{2}$ (27)
$+$ $\left(\frac{1}{2}-\frac{\sqrt{2}\sigma^{\prime}f_{\pi}+4\zeta^{\prime}(2f_{K}-f_{\pi})}{\sqrt{2}f^{2}}\right)\partial^{\mu}\eta_{8}\partial_{\mu}\eta_{8}$
$+\frac{d^{{}^{\prime}}}{4f^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)\partial^{\mu}\eta_{8}\partial_{\mu}\eta_{8},$
where $d^{\prime}$ = $3d_{1}+d_{2}$. Also, $\sigma^{{}^{\prime}}$ and $\zeta^{{}^{\prime}}$ are the fluctuations from the vacuum expectation values, i.e., $\sigma^{{}^{\prime}}=\sigma-\sigma_{0}$ and $\zeta^{{}^{\prime}}=\zeta-\zeta_{0}$. The first term of Eq.(27) (mass term) corresponds to the explicit symmetry breaking and is obtained from the general term
$\mathcal{L_{\text{Mass term}}}=-\frac{1}{2}{\rm Tr}A_{p}\left(uXu+u^{\dagger}Xu^{\dagger}\right).$ (28)
In the above equation, $A_{p}$ is a diagonal matrix given as
$A_{p}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}m_{\pi}^{2}f_{\pi}&0&0\\ 0&m_{\pi}^{2}f_{\pi}&0\\ 0&0&2m_{K}^{2}f_{K}-m_{\pi}^{2}f_{\pi}\end{array}\right).$ (29)
The vacuum mass $m_{\eta_{8}}$ of $\eta_{8}$ meson deduced from Eq.(28) is identified as
$m_{\eta_{8}}=\frac{1}{f}\sqrt{\frac{1}{2}(-\left(8f_{K}f_{\pi}(m_{\pi}^{2}+m_{K}^{2}))+16f_{K}^{2}m_{K}^{2}+6f_{\pi}^{2}m_{\pi}^{2}\right)},$ (30)
where $f=\sqrt{f_{\pi}^{2}+2\left(2f_{K}-f_{\pi}\right)^{2}}$. The second term of Eq.(27) is obtained from the general kinetic term of the pseudoscalar mesons (third term of Eq. (2)) and is also knwon as first range term in the chiral SU(3) model.
The last term of Eq.(27), in terms of $d_{1}$ and $d_{2}$, is obtained from the Lagrangian densities (known as $d_{1}$ and $d_{2}$ terms) [65, 66]
${\cal L}_{d_{1}}^{\bar{B}B}=\frac{d_{1}}{2}{\rm Tr}(u_{\mu}u^{\mu}){\rm Tr}(\bar{B}B),\quad\quad\quad\quad{\cal L}_{d_{2}}^{\bar{B}B}=d_{2}{\rm Tr}(\bar{B}u_{\mu}u^{\mu}B).$ (31)
In the present work for $\eta_{8}$ mesons, from the Weinberg Tomozawa term we do not obtain interaction terms with finite contributions unlike the case for kaons [65, 66] and $D$ [38] mesons studied in past using the present chiral SU(3) mean field model. These observations are also consistent with the calculations of $\eta$ mesons properties in the nuclear medium at zero temperature using chiral perturbation theory [80].
From the Lagrangian density given in Eq.(27), the equations of motion for $\eta_{8}$ is obtained, whose Fourier transformation leads to the dispersion relation
$-\omega^{2}+{\textbf{k}}^{2}+m_{\eta_{8}}^{2}+\Pi_{\eta_{8}}(\omega,|\textbf{k}|,\rho_{i}(r))=0,$ (32)
where $\Pi_{\eta_{8}}$ stands for $\eta_{8}$ meson’s self-energy and is given as,
$\Pi_{\eta_{8}}(\omega,|\textbf{k}|,\rho(r))$ $=$ $-\frac{8\zeta^{\prime}m^{2}_{K}f_{K}+(\sqrt{2}\sigma^{\prime}-4\zeta^{\prime})m^{2}_{\pi}f_{\pi}}{\sqrt{2}f^{2}}+\left[\frac{d^{\prime}}{2f^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)\right.$ (33)
-- $\frac{2}{f^{2}}\left(\left.\frac{\sqrt{2}\sigma^{\prime}f_{\pi}+4\zeta^{\prime}(2f_{K}-f_{\pi})}{\sqrt{2}}\right)\right](-\omega^{2}+{\textbf{ k}}^{2}).$
At zero momentum ($\textbf{k}=0$), the effective mass $m_{\eta_{8}}^{*}$ is written as
$m_{\eta_{8}}^{*}=\sqrt{m_{\eta_{8}}^{2}+\Pi_{\eta_{8}}(\omega,|\textbf{k}|,\rho_{i}(r))}.$ (34)
For the pseudoscalar singlet $\eta_{0}$, within the chiral SU(3) hadronic mean field model, we write
$\mathcal{L}=\frac{1}{2}\text{Tr}D_{\mu}YD^{\mu}Y-\frac{1}{2}m_{\eta_{0}}^{2}\,\text{Tr}Y^{2}-\frac{1}{2}\text{Tr}A_{p}\left(u(X+iY)u+u^{\dagger}(X-iY)u^{\dagger}\right).$ (35)
In above, the first term is the is the kinetic energy term for pseudoscalar mesons. The second term is the part of explicit symmetry breaking term and gives the finite mass to the pseudoscalar singlet. The last term of Eq.(35) is the extension of Eq.(28) where now the singlet $Y$ is included. This leads to the terms having mixing between $\eta_{0}$ and $\eta_{8}$ states. Explicitly, using Eq.(26) in Eq.(35), we get
$\mathcal{L}=\frac{1}{2}\partial_{\mu}\eta_{0}\partial^{\mu}\eta_{0}-\frac{1}{2}m_{\eta_{0}}^{2}\eta_{0}^{2}+\frac{1}{\sqrt{3}\sigma_{0}c}\left(m_{\pi}^{2}f_{\pi}-m_{K}^{2}f_{K}\right)\eta_{0}\eta_{8},$ (36)
where, $c={2\sigma_{0}}/{(\sqrt{2}\zeta_{0}+\sigma_{0})}$. In the last term of above equation the interaction terms corresponding to pure $\eta_{8}$ states, which are given in Eq.(27), are not repeated to avoid double counting. From Eq.(36) we observe that unlike $\eta_{8}$, for pure $\eta_{0}$ state contributions from the in-medium interactions does not appear, i.e.,
$m_{\eta_{0}}^{*}=m_{\eta_{0}}.$ (37)
From the mixed term (last term of Eq.(36)), we identify
$m_{\eta_{0}\eta_{8}}^{*}=\left[\frac{1}{\sqrt{3}\sigma_{0}c}\left(m_{\pi}^{2}f_{\pi}-m_{K}^{2}f_{K}\right)\right]^{1/2}.$ (38)
The effective masses $m_{\eta}^{*}$ and $m_{\eta^{{}^{\prime}}}^{*}$ of physical $\eta$ and $\eta^{{}^{\prime}}$ mesons are obtained by diagonalization of mass matrix
$m_{ij}=\begin{pmatrix}m_{\eta_{0}}^{*}&m_{\eta_{0}\eta_{8}}^{*}\\ m_{\eta_{8}\eta_{0}}^{*}&m_{\eta_{8}}^{*}\end{pmatrix}.$ (39)
The parameter $d^{{}^{\prime}}$ is expressed in terms of $\eta-N$ scattering length, $a^{\eta N}$. The expression for scattering length $a^{\eta N}$ calculated within chiral SU(3) model is given by [68]
$a^{\eta N}$ $=$ $\frac{1}{4\pi\left(1+\frac{m_{\eta}}{M_{N}}\right)}\Big{[}\Big{(}\frac{d^{\prime}}{\sqrt{2}}-\frac{g_{\sigma N}{f_{\pi}}}{m^{2}_{\sigma}}+\frac{4(2f_{K}-f_{\pi})g_{\zeta N}}{m^{2}_{\zeta}}\Big{)}\frac{m_{\eta}^{2}}{\sqrt{2}f^{2}}$ (40)
$+$ $\left(\frac{\sqrt{2}g_{\sigma N}}{m^{2}_{\sigma}}-\frac{4g_{\zeta N}}{m^{2}_{\zeta}}\right)\frac{m^{2}_{\pi}f_{\pi}}{2\sqrt{2}f^{2}}+\frac{2\sqrt{2}g_{\delta N}}{m^{2}_{\delta}}\frac{m^{2}_{K}f_{K}}{f^{2}}\Big{]},$
which can be rearranged to obtain the expression for $d^{{}^{\prime}}$. We calculate $d^{\prime}$ for the value of $a^{\eta N}=$ 0.91 [80]. From the above calculated effective masses $m_{\eta}^{*}$ and $m_{\eta^{\prime}}^{*}$ of $\eta$ and $\eta^{\prime}$ mesons, corresponding mass shift will be calculated which shall be used as input in Eqs. (22) and (24) to obtain the optical potentials of these mesons in the nuclei.
III.2 Pseudoscalar $D^{0}$ and $\bar{D^{0}}$ mesons
The effective masses of neutral $D^{0}$ and $\bar{D^{0}}$ mesons required in Eq.(22) to study the $D$ mesic nuclei are obtained in the present work by extending the chiral SU(3) model to the SU(4) case [38]. The mesons $D^{0}$ and $\bar{D^{0}}$ are the members of $D$ $\left(D^{+},D^{0}\right)$ and $\bar{D}$ $\left(D^{-},\bar{D^{0}}\right)$ doublets, respectively. From the Lagrangian density describing the interactions of $D$ and $\bar{D}$ mesons with nucleons in the nuclear medium, equations of motion are obtained whose Fourier transformation will give the dispersion relation
$-\omega^{2}+{\textbf{k}}^{2}+m_{D^{0},\bar{D^{0}}}^{2}+\Pi_{D^{0},\bar{D^{0}}}(\omega,|\textbf{k}|,\rho_{i}(r))=0.$ (41)
For the $D^{0}$ mesons, the expression of self-energy is given by
$\Pi_{D^{0}}(\omega,|\textbf{k}|,\rho(r))$ $=\frac{1}{2f_{D}^{2}}\left(2\rho_{p}+\rho_{n}\right)\omega+\frac{m_{D}^{2}}{2f_{D}}\left(\sigma^{\prime}+\sqrt{2}\zeta_{c}^{\prime}+\delta^{\prime}\right)$
$+\left[-\frac{1}{f_{D}}\left(\sigma^{\prime}+\sqrt{2}\zeta_{c}^{\prime}+\delta^{\prime}\right)+\frac{d_{1}}{2f_{D}^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)\right.\left.+\frac{d_{2}}{2f_{D}^{2}}\rho_{p}^{s}\right]\left(\omega^{2}-\textbf{k}^{2}\right).$ (42)
For $\bar{D^{0}}$, we have
$\Pi_{\bar{D^{0}}}(\omega,|\textbf{k}|,\rho(r))$ $=-\frac{1}{2f_{D}^{2}}\left(2\rho_{p}+\rho_{n}\right)\omega+\frac{m_{D}^{2}}{2f_{D}}\left(\sigma^{\prime}+\sqrt{2}\zeta_{c}^{\prime}+\delta^{\prime}\right)$
$+\left[-\frac{1}{f_{D}}\left(\sigma^{\prime}+\sqrt{2}\zeta_{c}^{\prime}+\delta^{\prime}\right)+\frac{d_{1}}{2f_{D}^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)+\frac{d_{2}}{2f_{D}^{2}}\rho_{p}^{s}\right]\left(\omega^{2}-\textbf{k}^{2}\right).$ (43)
In Eqs. (42) and (43), $\zeta_{c}^{\prime}$ denote the fluctuations of charm condensate $(c\bar{c})$ from the vacuum expectation value and being heavy in flavor, is considered zero in the current work. From the dispersion relation, the effective mass for the $D$ mesons (at zero momentum) is written as
$m_{D^{0},\bar{D^{0}}}^{*}=\sqrt{m_{D^{0},\bar{D^{0}}}^{2}+\Pi_{D^{0},\bar{D^{0}}}(\omega,|\textbf{k}|,\rho_{i}(r))}$ (44)
The in-medium mass of $D$ mesons calculated using the above relation is used in Eqs. (22) and (24), to obtain the real and imaginary optical potentials.
III.3 Pseudoscalar $B^{0}$ and $\bar{B^{0}}$ mesons
To study the formation of bound states corresponding to $B^{0}$ and $\bar{B^{0}}$ mesons chiral SU(3) model will be generalized to SU(5) sector as these bottom mesons are composed of one heavy $b$ quark/antiquark. The pseudoscalar $B^{0}$ and $\bar{B^{0}}$ mesons belong to the open bottom $B(B^{+},B^{0})$ and $\bar{B}(B^{-},\bar{B^{0}})$ meson doublets. In Ref. [43] the masses of open bottom mesons have been calculated in nuclear and strange hadronic matter using chiral effective model. For our present calculations, we need the mass shift of $B^{0}$ and $\bar{B^{0}}$ mesons in the nuclear medium at zero temperature obtained by solving the dispersion relation
$-\omega^{2}+{\textbf{k}}^{2}+m_{B^{0},\bar{B^{0}}}^{2}+\Pi_{B^{0},\bar{B^{0}}}(\omega,|\textbf{k}|,\rho_{i}(r))=0.$ (45)
In case of $B^{0}$ and $\bar{B^{0}}$ mesons, the expressions of self-energies are given by [43]
$\Pi_{B^{0}}(\omega,|\textbf{k}|,\rho(r))$ $=-\frac{1}{2f_{B}^{2}}\left(\rho_{p}+2\rho_{n}\right)\omega+\frac{m_{B}^{2}}{2f_{B}}\left(\sigma^{\prime}+\sqrt{2}\zeta_{b}^{\prime}-\delta^{\prime}\right)$
$+\left[-\frac{1}{f_{B}}\left(\sigma^{\prime}+\sqrt{2}\zeta_{b}^{\prime}+\delta^{\prime}\right)+\frac{d_{1}}{2f_{B}^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)\right.\left.+\frac{d_{2}}{2f_{B}^{2}}\left(\rho_{p}^{s}+2\rho_{n}^{s}\right)\right]\left(\omega^{2}-\textbf{k}^{2}\right),$ (46)
and
$\Pi_{\bar{B^{0}}}(\omega,|\textbf{k}|,\rho(r))$ $=\frac{1}{2f_{B}^{2}}\left(\rho_{p}+2\rho_{n}\right)\omega+\frac{m_{B}^{2}}{2f_{B}}\left(\sigma^{\prime}+\sqrt{2}\zeta_{b}^{\prime}-\delta^{\prime}\right)$
$+\left[-\frac{1}{f_{B}}\left(\sigma^{\prime}+\sqrt{2}\zeta_{b}^{\prime}+\delta^{\prime}\right)+\frac{d_{1}}{2f_{B}^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)\right.\left.+\frac{d_{2}}{2f_{B}^{2}}\left(\rho_{p}^{s}+2\rho_{n}^{s}\right)\right]\left(\omega^{2}-\textbf{k}^{2}\right),$ (47)
respectively. Similar to $\zeta_{c}^{\prime}$, the fluctuations $\zeta_{b}^{\prime}$ (corresponding the condensate, $b\bar{b}$, of heavy $b$ quark) appearing in Eqs. (46) and (47), are considered zero in the calculations of mass shift of $B^{0}$ and $\bar{B^{0}}$ mesons. At zero momentum, the in-medium masses of neutral $B$ mesons required to obtain the mass shift and hence, the optical potentials in different nuclei, can be written as
$m_{B^{0},\bar{B^{0}}}^{*}=\sqrt{m_{B^{0},\bar{B^{0}}}^{2}+\Pi_{B^{0},\bar{B^{0}}}(\omega,|\textbf{k}|,\rho_{i}(r))}.$ (48)
III.4 Pseudoscalar $\bar{K^{0}}$ mesons
The in-medium masses of kaons and antikaons have been studied using chiral SU(3) model in nuclear [65, 66] and strange hadronic medium [64]. In the nuclear matter, the effective mass of kaons $(K^{+},K^{0})$ is observed to increase whereas the mass of antikaons $(K^{-},\bar{K^{0}})$ decreases as a function of density of the nuclear medium. Since the negative mass shift implies the attractive optical potentials are required to form the bound states with nuclei, here we study the possible formation of bound states of $\bar{K^{0}}$ mesons with light and heavy nuclei. Similar to Eqs. (41) and (45), within the chiral model, the effective mass of $\bar{K^{0}}$ mesons will be obtained by solving the dispersion relation, for which the required self energy $\Pi_{\bar{K^{0}}}(\omega,|\textbf{k}|,\rho(r))$ is written as [66]
$\Pi_{\bar{K^{0}}}(\omega,|\textbf{k}|,\rho(r))$ $=\frac{1}{2f_{K}^{2}}\left(\rho_{p}+2\rho_{n}\right)\omega+\frac{m_{B}^{2}}{2f_{B}}\left(\sigma^{\prime}+\sqrt{2}\zeta^{\prime}-\delta^{\prime}\right)$
$+\left[-\frac{1}{f_{K}}\left(\sigma^{\prime}+\sqrt{2}\zeta^{\prime}+\delta^{\prime}\right)+\frac{d_{1}}{2f_{K}^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)\right.\left.+\frac{d_{2}}{2f_{K}^{2}}\rho_{n}^{s}\right]\left(\omega^{2}-\textbf{k}^{2}\right).$ (49)
III.5 Vector $\phi$ mesons
To investigate the possibility of the formation of $\phi$ mesic nuclei, we calculate the effective masses and decay width of $\phi$ mesons using the effective Lagrangian which takes into account the interactions of $\phi$ mesons with kaons, $K\left(K^{+},K^{0}\right)$ and antikaons, $\bar{K}\left(K^{-},\bar{K^{0}}\right)$ isospin doublets. This interaction Lagrangian density is written as [81, 82]
$\mathcal{L}_{\text{int }}=ig_{\phi}\phi^{\mu}\left[\bar{K}\left(\partial_{\mu}K\right)-\left(\partial_{\mu}\bar{K}\right)K\right].$ (50)
Using above, $\phi$ meson self-energy, $\Pi_{\phi}^{*}(p)$, in the nuclear medium relevant for different nuclei, is calculated for the decay process $\phi\rightarrow K\bar{K}$ at the one-loop In Eq. (50), $g_{\phi}$ represents the coupling constant. Since the contributions of $\phi\phi$$\bar{K}$ interactions to the in-medium masses and decay width are smaller than those of $\phi$$\bar{K}$ interactions, we have not taken into account these interactions in our current work on the interaction Lagrangian [82]. The in-medium mass of the $\phi$ meson is expressed in terms of real part of self energy, $\Pi_{\phi}^{*}(p)$, through the relation
$m_{\phi}^{*^{2}}=\left(m_{\phi}^{0}\right)^{2}+\operatorname{Re}\Pi_{\phi}^{*}\left(m_{\phi}^{*^{2}}\right).$ (51)
In the above equation, $m_{\phi}^{0}$ is the bare mass of the $\phi$ meson. Also, the real part of the self energy is expressed in terms of in-medium energies of kaons and antikaons as [59, 82]
$\operatorname{Re}\Pi_{\phi}^{*}=-\frac{4}{3}g_{\phi}^{2}\mathcal{P}\int\frac{d^{3}q}{(2\pi)^{3}}\vec{q}^{2}\frac{\left(E_{K}^{*}+E_{\bar{K}}^{*}\right)}{E_{K}^{*}E_{\bar{K}}^{*}\left(\left(E_{K}^{*}+E_{\bar{K}}^{*}\right)^{2}-m_{\phi}^{*^{2}}\right)},$ (52)
[…]
$\operatorname{Re}\Pi_{\phi}^{*}=-\frac{4}{3}g_{\phi}^{2}\mathcal{P}\int_{0}^{\Lambda_{c}}\frac{d^{3}q}{(2\pi)^{3}}\vec{q}^{2}\left(\frac{\Lambda_{c}^{2}+m_{\phi}^{*^{2}}}{\Lambda_{c}^{2}+4E_{K}^{*^{2}}}\right)^{4}\frac{\left(E_{K}^{*}+E_{\bar{K}}^{*}\right)}{E_{K}^{*}E_{\bar{K}}^{*}\left(\left(E_{K}^{*}+E_{\bar{K}}^{*}\right)^{2}-m_{\phi}^{*^{2}}\right)}.$ (53)
For the case of $\phi$ meson, the value of decay width required in Eq. (23), is obtained from the imaginary component of its self energy. The expression of decay width, expressed in terms of in-medium masses of $\phi$ mesons as well as the masses of kaons and antikaons, is given by [84]
$\Gamma_{\phi}^{*}=\frac{g_{\phi}^{2}}{24\pi}\frac{1}{m_{\phi}^{*^{5}}}\left[\left(m_{\phi}^{*^{2}}-\left(m_{K}^{*}+m_{\bar{K}}^{*}\right)^{2}\right)\right.\left.\times\left(m_{\phi}^{*^{2}}-\left(m_{K}^{*}-m_{\bar{K}}^{*}\right)^{2}\right)\right]^{3/2}.$ (54)
The value of coupling constant $g_{\phi}$ is fixed to empirical width of the $\phi$ meson in the vacuum and found to be $4.539$. Also, the bare mass of the $\phi$ meson appearing in Eq. (51) is calculated by setting its vacuum mass at 1019.461 MeV.
IV Results and discussion
We now discuss the possibility of the formation of mesic nuclei of various mesons through the numerical calculations of the optical potentials and binding energies. We shall calculate the optical potentials of $\eta,\eta^{{}^{\prime}}$, $D^{0}$, $\bar{D^{0}}$, $B^{0}$, $\bar{B^{0}}$,$\bar{K^{0}}$ and $\phi$ mesons in the medium relevant for the nuclei 12C, 16O, 40Ca and 208Pb. As we discussed in the previous section, the in-medium properties of mesons are calculated using the chiral SU(3) hadronic mean field model and extended to SU(4) and SU(5) sector to obtain the in-medium masses of charmed and bottom mesons [63].
$A$ $nl$ $\cal E_{B}$ (MeV) $\Gamma$ (MeV)
$\kappa$ =0 $\kappa$ = 0.5 $\kappa$ =1 $\kappa$ =0 $\kappa$ = 0.5 $\kappa$ =1
${}^{12}_{6}$ $1s$0.511 -0.103 0.0 2.142
${}^{16}_{8}$ $1s$1.72 -1.38 -0.458 0.0 3.775 8.094
${}^{40}_{20}$a $1s$9.769 -9.538 -8.91 0.0 9.527 19.387
${}^{208}_{82}$b $1s$20.642 -20.555 - 20.309 0.0 12.987 26.046
$1p$14.444 -14.303 -13.921 0.0 11.996 24.175
$1d$7.244 -7.005 -6.392 0.0 10.621 21.656
$2s$5.123 -4.74 -3.846 0.0 9.544 19.896
Table 1: Values of binding energies, $\cal E_{B}$, and full decay width, $\Gamma$, for $\eta$ mesons in four nuclei with different mass number $A$ are tabulated for $\kappa=0,0.5$ and 1.
Figure 1: The mass shift of $\eta$ [in subplot (a)] and $\eta^{\prime}$ [in subplot (b)] mesons are shown as a function of density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$.
Figure 2: The real optical potential $U(r)$ of $\eta$ [in subplot (a)] and $\eta^{\prime}$ [in subplot (b)] mesons are shown as a function of distance $r$ from center, for nuclei 12C, 16O, 40Ca and 208Pb, with mass number, $A=12,16,40$ and $208$, respectively.
IV.1 Binding energies for $\eta$ and $\eta^{{}^{\prime}}$ mesic nuclei
In this subsection, we shall explore the possibility of the formation of $\eta$ and $\eta^{{}^{\prime}}$ mesic nuclei. Using Eq. (39), the mass shifts of pseudoscalar $\eta$ and $\eta^{{}^{\prime}}$ mesons are calculated within the chiral SU(3) model. In Fig. 1, we have shown the mass shift of $\eta$ and $\eta^{{}^{\prime}}$ mesons as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of the infinite nuclear medium for isospin asymmetry $I=0,0.211$ and $0.3$. The isospin asymmetry $I=0.211$ corresponds to the heavy nuclei 208Pb considered in the present work for the bound state formation. As we discussed earlier, for the calculations of $\eta$ and $\eta^{{}^{\prime}}$ meson properties, we considered the mixing of $\eta_{8}$ and $\eta_{0}$ states. The pseudoscalar $\eta$ meson which is member of octet is observed to undergo larger mass shift as compared to the singlet $\eta^{{}^{\prime}}$. At nuclear saturation density $\rho_{B}=\rho_{0}$, for isospin asymmetry $I=0(0.3)$, the values of mass shift $\Delta m_{\eta}$ and $\Delta m_{\eta}^{{}^{\prime}}$ are observed to be $-34.64(-34.21)$ and $-1.03(-1.02)$ MeV, respectively. At baryon density $\rho_{B}=4\rho_{0}$, the above values of mass shift change to $-124.15(-120.92)$ and $-3.00(-2.94)$ MeV. The finite isopsin asymmetry of the medium causes less drop in the masses of $\eta$ and $\eta^{{}^{\prime}}$, although the impact is small. Furthermore, in the present chiral model calculations, compared to $\eta$ mesons, the singlet $\eta^{{}^{\prime}}$ undergoes very less drop in mass as a function of density of nuclear medium.
To study the bound states, the Klein Gordon equation [Eq. (20)], which require the optical potential corresponding to different nuclei under study as input is solved. In Fig. 2 the real part of the optical potentials, for $\eta$ and $\eta^{\prime}$ mesons, calculated using Eq. (22) is plotted as a function of distance $r$ from the center of nuclei. The heavy nuclei are observed to have larger value of negative optical potential values as compared to lighter nuclei. The imaginary optical potential which corresponds to the absorption of mesons in the nuclei is calculated using Eq. (23). Using the real and imaginary values of optical potential in the Klein Gordon equation, the binding energy and absorption decay width are calculated. In case of $\eta$ mesons, the values of binding energies and decay width are tabulated in Table 1. In case of light nuclei 12C the negative values for the binding energies are observed only in $1s$ state for the value of phenomenological parameter $\kappa=0$ and $0.5$. As described earlier in Sec. III, through the parameter $\kappa$ we simulate the impact of imaginary optical potential on the formation of bound states. As the value of $\kappa$ is increased to $1$, no bound state is observed. For the case of 16O and 40Ca, bound state formation in the $1s$ state may take place for all values of $\kappa$ considered in the present work. As can be seen from Table 1, in the case of 208Pb nuclei the negative values for the binding energies are observed not only for the $1s$ state, but also for the other excited states, for example, $1p,1d$ and $2s$ as well. Due to larger value of attractive optical potential (magnitude) of $\eta$ meson with heavy 208Pb nuclei, the possibility of the formation of bound state is more with these nuclei.
As discussed earlier, in the present calculations the $\eta^{{}^{\prime}}$ mesons are not observed to undergo appreciable mass drop in the nuclear medium ($\Delta m_{\eta^{{}^{\prime}}}=-1.03$ MeV at $\rho_{B}=\rho_{0}$) and bound states are not observed for these pseudoscalar singlet meson with any of the nuclei considered in the present work. The present results of $\eta$ and $\eta^{{}^{\prime}}$ mesons can be compared with the earlier studies using different approaches. In Ref. [14] the mass shifts of $\eta$ and $\eta^{{}^{\prime}}$ mesons were calculated in the symmetric nuclear medium using the QMC model considering the $\eta-\eta^{{}^{\prime}}$ mixing. The values of mass shift are observed to be $-63.6$ and $-61.3$ MeV, respectively, at $\rho_{B}=\rho_{0}$. The binding energies and absorption decay width were calculated for light and heavy nuclei and were observed to form the bound states due to significant mass drop. Using the coupled channel approach, the optical potential for $\eta^{{}^{\prime}}$ mesons was found to be $-(8.7+1.8i)$ MeV at nuclear saturation density, $\rho_{0}$ and no bound state is observed in Ref. [9]. Bound states of $\eta$ mesons have also been explored using unitarized coupled channel approach in Ref. [17], considering the energy independent and dependent potentials. The binding energies obtained were larger compared to our present calculations. The feasibility of the formation of $\eta$ and $\eta^{{}^{\prime}}$ meson bound states with nuclei has also been explored using the NJL model [8]. In Ref. [85], relativistic field theoretical model under mean field approximation is used to study the $\eta^{\prime}$ mesic nuclei through the direct coupling of $\eta^{\prime}$ meson with scalar mean field $\sigma$ (interaction of form $\sim g_{\sigma\eta^{\prime}}m_{\eta^{\prime}}\eta^{\prime 2}\sigma$). Assuming the partial restoration of chiral symmetry and $30$% reduction in the chiral condensate to determine the coupling constant $g_{\sigma\eta^{\prime}}$, the $\eta^{\prime}$ meson is observed to undergo a mass drop of $80$ MeV at nuclear saturation density from vacuum value. Such larger mass drop obviously lead to the formation of bound states with nuclei 12C, 16O and 40Ca, considered in the work.
$A$ $nl$ $\cal E_{B}$ (MeV) $\Gamma$ (MeV)
$\kappa$ =0 $\kappa$ = 0.5 $\kappa$ =1 $\kappa$ =0 $\kappa$ = 0.5 $\kappa$ =1
${}^{12}_{6}$ 1s -25.965 -25.741 -25.118 0.0 18.731 37.748
1p -11.246 -10.811 -9.638 0.0 14.235 29.144
2s -0.935 0.0
${}^{16}_{8}$ 1s -29.11 -28.941 -28.474 0.0 19.504 39.224
1p -16.516 -16.184 -15.282 0.0 16.451 33.407
1d -3.884 -3.234 -1.619 0.0 12.057 25.31
2s -3.946 -3.152 -1.182 0.0 9.717 20.971
${}^{40}_{20}$a 1s -47.743 -47.607 -47.22 0.0 27.865 55.842
1p -38.223 -38.016 -37.44 0.0 26.065 52.369
1d -27.642 -27.338 -26.509 0.0 23.945 48.313
2s -25.594 -25.233 -24.263 0.0 23.249 47.047
2p -14.212 -13.608 -12.07 0.0 19.93 40.966
2d -3.852 -2.7 -0.152 0.0 15.302 33.133
${}^{208}_{82}$b 1s -49.133 -49.073 -48.897 0.0 26.019 52.054
1p -46.417 -46.344 -46.135 0.0 25.734 51.507
1d -43.096 -43.008 -42.757 0.0 25.382 50.832
2s -41.77 -41.674 -41.401 0.0 25.235 50.554
2p -37.132 -37.008 -36.664 0.0 24.711 49.56
2d -32.078 -31.919 -31.486 0.0 24.102 48.414
Table 2: Values of binding energies, $\cal E_{B}$, and full decay width, $\Gamma$ for $D^{0}$ mesons in four nuclei with different mass number $A$ are tabulated for $\kappa=0,0.5$ and 1.
$A$ $nl$ $\cal E_{B}$ (MeV) $\Gamma$ (MeV)
$\kappa$ =0 $\kappa$ = 0.5 $\kappa$ =1 $\kappa$ =0 $\kappa$ = 0.5 $\kappa$ =1
${}^{12}_{6}$ 1s -1.633 -1.49 -1.102 0.0 2.539 5.306
${}^{16}_{8}$ 1s -2.142 -2.03 -1.726 0.0 2.756 5.694
${}^{40}_{20}$a 1s -9.87 -9.801 -9.615 0.0 7.096 14.299
1p -4.25 -4.091 -3.676 0.0 5.674 11.637
${}^{208}_{82}$b 1s -13.541 -13.524 -13.479 0.0 7.646 15.313
1p -11.434 -11.403 -11.32 0.0 7.41 14.869
1d -8.887 -8.836 -8.701 0.0 7.101 14.29
2s -7.936 -7.869 -7.694 0.0 6.924 13.973
2p -4.607 -4.478 -4.164 0.0 6.303 12.881
2d -1.3 -1.007 -0.43 0.0 5.311 11.331
Table 3: Values of binding energies, $\cal E_{B}$, and full decay width, $\Gamma$, for $\bar{D^{0}}$ mesons in four nuclei with different mass number $A$ are tabulated for $\kappa=0,0.5$ and 1.
IV.2 Binding energies for $D^{0}$ and $\bar{D^{0}}$ mesic nuclei
Now we discuss the formation of bound states for neutral $D^{0}$ and $\bar{D^{0}}$ mesons with the nuclei within the chiral SU(3) hadronic mean field model. The mass shift of $D^{0}$ and $\bar{D^{0}}$ mesons calculated using Eq. (44) are plotted in Fig. (3) as a function of baryon density ratio $\rho_{B}/\rho_{0}$. As can be seen from the figure, the in-medium masses of both $D^{0}$ and $\bar{D^{0}}$ mesons decreases as a function of density of the nuclear medium. Furthermore, the increase of isospin asymmetry of the medium from zero to finite value (which is relevant for the study of nuclei having different number of protons and neutrons) decreases the magnitude of mass shift for these neutral pseudoscalar mesons. At baryon density $\rho_{B}=\rho_{0}(4\rho_{0})$, the values of mass shift for $D^{0}$ and $\bar{D^{0}}$ mesons are observed to $-76.96(-343.82)$ and $-26(-159.71)$ MeV in the symmetric nuclear medium, whereas for $\eta=0.3$ these values changes to $-64.17(-298.80)$ and $-23.37(-148.43)$ MeV, respectively. Using the observed values of mass shift for $D^{0}$ and $\bar{D^{0}}$ mesons at center of given nuclei in Eq. (22), the real part of optical potential is calculated and is plotted in Fig. 4 as a function of $r$ for nuclei, 12C, 16O, 40Ca and 208Pb. As one can see from the figure, depth of optical potential (large negative value) is observed to be more for $D^{0}$ mesons as compared to $\bar{D^{0}}$. Furthermore, the imaginary part of optical potential, required to understand the absorption of mesons in the nuclei and its impact on the binding energy of neutral $D$ mesic nuclei, is calculated using Eq. (23). The values of binding energy and decay width obtained by solving the Klein Gordon equation are given in Tables 2 and 3, for $D^{0}$ and $\bar{D^{0}}$ mesons, respectively, for $\kappa=0,0.5$ and 1. As one can see from Table 2, the bound states of $D^{0}$ mesons, for 40Ca and 208Pb may be found not only for lower states $1s$ and $1p$, but also for other higher states when $\kappa=0$ and $0.5$. When the value of $\kappa$ is increased to one, the magnitude of binding energy is observed to be less as compared to the decay width value. This may cause a broadening of the peak and hindrance in the observation of clear signature of bound states. Comparing the values of binding energies in Tables 2 and 3, magnitude of binding energies for $\bar{D^{0}}$ mesons are significantly less than $D^{0}$ mesons. In Ref. [10], the bound states of $D^{0}$ mesons have been studied for several nuclei using the coupled channel approach. Comparing our results for 12C, 40Ca and 208Pb with Ref. [10], the magnitude of binding energies are observed to be more in our case. The binding energies of $D^{0}$ and $\bar{D^{0}}$ mesons for 208Pb nuclei are calculated using QMC model and are found to approximately $-96$ and $-25$ MeV, respectively [7].
Figure 3: The mass shift of $D^{0}$ [in subplot (a)] and $\bar{D^{0}}$ [in subplot (b)] mesons are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$.
Figure 4: The real optical potential $U(r)$ of $D^{0}$ [in subplot (a)] and $\bar{D^{0}}$ [in subplot (b)] mesons are shown as a function of distance $r$ from center, for nuclei 12C, 16O, 40Ca and 208Pb, with mass number, $A=12,16,40$ and $208$, respectively.
Figure 5: The mass shift of $B^{0}$ [in subplot (a)] and $\bar{B^{0}}$ [in subplot (b)] mesons are shown as a function of density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$.
Figure 6: The real optical potential $U(r)$ of $B^{0}$ [in subplot (a)] and $\bar{B^{0}}$ [in subplot (b)] mesons are shown as a function of distance $r$ from center, for nuclei 12C, 16O, 40Ca and 208Pb, with mass number, $A=12,16,40$ and $208$, respectively.
IV.3 Binding energies for $B^{0}$ and $\bar{B^{0}}$ mesic nuclei
The studies on the bound states of $B$ meson with nuclei has not been done much in literature [44]. In the present section we shall discuss the formation of bound states of pseudoscalar $B^{0}$ and $\bar{B^{0}}$ mesons by calculating their optical potential using the chiral SU(3) model. In Fig. 5, the mass shift of these pseudoscalar mesons are plotted as a function of baryon density $\rho_{B}$ (in units of $\rho_{0}$) for isospin asymmetry $I=0,0.211$ and $0.3$. Both pseudoscalar $B^{0}$ and $\bar{B^{0}}$ mesons are observed to undergo appreciable mass drop at nuclear matter density. However, the isospin asymmetry effects are not appreciable. In symmetric nuclear medium the values of $\Delta m$ for $B^{0}$ and $\bar{B^{0}}$ mesons are observed to be $-48.69$ and $-74.06$ MeV, whereas at $4\rho_{0}$ these values changes to $-356.40$ and $-453.28$ MeV, respectively. As shown in Fig. 6, for a given nuclei, at the center, the real part of optical potentials have larger value for $\bar{B^{0}}$ as compared to $B^{0}$ mesons. The values of binding energy and decay width for these pseudoscalar mesosn are given in Table 4 and 5. As expected from the magnitude of mass drop for $B^{0}$ and $\bar{B^{0}}$ mesons, the magnitude of binding energy with given nuclei is more for $\bar{B^{0}}$ mesons as compared to $B^{0}$.
$A$ $nl$ $\cal E_{B}$ (MeV) $\Gamma$ (MeV)
$\kappa$ =0 $\kappa$ = 0.5 $\kappa$ =1 $\kappa$ =0 $\kappa$ = 0.5 $\kappa$ =1
${}^{12}_{6}$ 1s -10.933 -10.845 -10.604 0.0 7.907 15.944
1p -4.572 -4.385 -3.886 0.0 5.989 12.285
2s -0.272 0.0
${}^{16}_{8}$ 1s -10.848 -10.789 -10.627 0.0 7.362 14.816
1p -5.789 -5.657 -5.302 0.0 6.133 12.486
1d -0.805 -0.513 0.0 4.274
2s -1.02 -0.662 0.0 3.274
${}^{40}_{20}$a 1s -26.07 -26.032 -25.926 0.0 14.613 29.275
1p -22.064 -21.996 -21.809 0.0 13.928 27.954
1d -17.562 -17.456 -17.166 0.0 13.13 26.425
2s -16.627 -16.502 -16.166 0.0 12.902 26.009
2p -11.516 -11.316 -10.79 0.0 11.741 23.84
2d -6.481 -6.158 -5.352 0.0 10.282 21.205
${}^{208}_{82}$b 1s -31.71 -31.698 -31.665 0.0 16.342 32.691
1p -30.648 -30.632 -30.586 0.0 16.245 32.505
1d -29.347 -29.325 -29.263 0.0 16.125 32.277
2s -28.826 -28.802 -28.733 0.0 16.08 32.19
2p -26.996 -26.961 -26.868 0.0 15.91 31.867
2d -24.983 -24.938 -24.815 0.0 15.717 31.502
Table 4: Values of binding energies, $\cal E_{B}$, and full decay width, $\Gamma$ for $B^{0}$ mesons in four nuclei with different mass number $A$ are tabulated for $\kappa=0,0.5$ and 1.
$A$ $nl$ $\cal E_{B}$ (MeV) $\Gamma$ (MeV)
$\kappa$ =0 $\kappa$ = 0.5 $\kappa$ =1 $\kappa$ =0 $\kappa$ = 0.5 $\kappa$ =1
${}^{12}_{6}$ 1s -25.403 -25.292 -24.986 0.0 16.151 32.455
1p -16.432 -16.222 -15.646 0.0 13.936 28.186
1d -7.473 -7.117 -6.171 0.0 11.205 22.994
2s -7.506 -7.101 -6.028 0.0 10.381 21.446
2p -1.086 -0.26 0.0 6.142
${}^{16}_{8}$ 1s -26.085 -26.014 -25.818 0.0 15.715 31.532
1p -19.262 -19.116 -18.717 0.0 14.435 29.096
1d -11.721 -11.471 -10.799 0.0 12.632 25.674
2s -10.487 -10.193 -9.407 0.0 11.52 23.535
2p -4.153 -3.645 -2.362 0.0 8.743 18.447
${}^{40}_{20}$a 1s -45.779 -45.725 -45.572 0.0 24.844 49.743
1p -41.003 -40.917 -40.677 0.0 24.034 48.174
1d -35.617 -35.492 -35.146 0.0 23.119 46.409
2s -34.457 -34.317 -33.934 0.0 22.917 46.034
2p -28.121 -27.919 -27.371 0.0 21.716 43.749
2d -21.612 -21.326 -20.57 0.0 20.34 41.161
${}^{208}_{82}$b 1s -55.277 -55.252 -55.179 0.0 28.274 56.555
1p -54.102 -54.073 -53.986 0.0 28.164 56.343
1d -52.663 -52.627 -52.523 0.0 28.029 56.085
2s -52.09 -52.052 -51.941 0.0 27.981 55.992
2p -50.065 -50.017 -49.883 0.0 27.796 55.639
2d -47.833 -47.775 -47.612 0.0 27.59 55.244
Table 5: Values of binding energies, $\cal E_{B}$, and full decay width, $\Gamma$ for $\bar{B^{0}}$ mesons in four nuclei with different mass number $A$ are tabulated for $\kappa=0,0.5$ and 1.
$A$ $nl$ $\cal E_{B}$ (MeV) $\Gamma$ (MeV)
$\kappa$ =0 $\kappa$ = 0.5 $\kappa$ =1 $\kappa$ =0 $\kappa$ = 0.5 $\kappa$ =1
${}^{12}_{6}$ 1s -0.573 -0.128 0.0 2.368
${}^{16}_{8}$ 1s -1.84 -1.469 -0.462 0.0 4.071 8.731
${}^{40}_{20}$a 1s -11.209 -10.949 -10.24 0.0 10.803 21.964
${}^{208}_{82}$b 1s -28.644 -28.509 -28.123 0.0 17.663 35.389
1p -21.387 -21.202 -20.688 0.0 16.485 33.146
1d -12.892 -12.621 -11.9 0.0 14.939 30.257
2s -10.123 -9.746 -8.799 0.0 13.919 28.501
2p -0.991 0.0
Table 6: Values of binding energies, $\cal E_{B}$, and full decay width, $\Gamma$ for $\bar{K^{0}}$ mesons in four nuclei with different mass number $A$ are tabulated for $\kappa=0,0.5$ and 1.
Figure 7: The mass shift of $\bar{K^{0}}$ mesons is shown as a function of density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$.
Figure 8: The real optical potential $U(r)$ of $\bar{K^{0}}$ mesons is shown as a function of distance $r$ from center, for nuclei 12C, 16O, 40Ca and 208Pb, with mass number, $A=12,16,40$ and $208$, respectively.
Figure 9: The mass shift [in subplot (a)] and decay width [in subplot (b)] of $\phi$ mesons are shown as a function of density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$.
Figure 10: The real optical potential $U(r)$ [in subplot (a)] and imaginary optical potential $W(r)$ [in subplot (b)] of $\phi$ mesons are shown as a function of distance $r$ from center, for nuclei 12C, 16O, 40Ca and 208Pb, with mass number, $A=12,16,,40$ and $208$, respectively.
IV.4 Binding energies for $\bar{K^{0}}$ and $\phi$ mesic nuclei
Lastly, in this subsection we explore the formation of bound states for $\bar{K^{0}}$ and $\phi$ mesons. Within the chiral SU(3) model, among the neutral $K^{0}$ and $\bar{K^{0}}$ mesons, only latter is observed to undergo mass drop as a function of density of the nuclear medium [65, 66, 64] and therefore, we consider $\bar{K^{0}}$ neutral meson only to explore the bound state formation. In Fig. 7 we have shown the variation of mass shift for $\bar{K^{0}}$ mesons as a function of baryon density $\rho_{B}$. The plot of real part of optical potential $U(r)$ as a function of distance $r$ from the center of the given nuclei of mass number $A$ is given for $\bar{K^{0}}$ meson in Fig. 8. At $\eta=0$, the values of mass shift for $\bar{K^{0}}$ mesons are found to be $-36.90$ and $-184.31$ MeV, at $\rho_{B}=\rho_{0}$ and $4\rho_{0}$, respectively. For a given density of medium, an increase in the isospin asymmetry enhances the mass drop for $\bar{K^{0}}$ mesons, which is more appreciable at higher baryon densities. As can be seen from Table 6, for the nuclei 12C, 18O and 40Ca, the negative values of binding energies are observed only for $1s$ state. Further, in this case when the value of $\kappa$ is considered finite, it may be difficult to observe the bound state of $\bar{K^{0}}$ meson with these nuclei due to comparable or larger value of decay width $\Gamma$.
As we discussed earlier in Sec. III.5, the mass shift and decay width of $\phi$ meson are obtained in the present work by calculating the self energy, which involve the interactions of $\phi$ mesons with $K$ and $\bar{K}$, at one loop level. In Fig. 9 we have plotted the mass shift and decay width of $\phi$ mesons as a function of $\rho_{B}/\rho_{0}$. The values of mass shift $\Delta m_{\phi}^{*}$ and decay width $\ \Gamma_{\phi}^{*}$ at baryon density $\rho_{B}=\rho_{0}(4\rho_{0})$ are observed to be $-2.00(-20.28)$ and $4.96(27.68)$ MeV, respectively, for the value of cuttoff parameter, $\Lambda_{c}=3000$ MeV. In Fig. 10 the real and imaginary optical potentials, $U_{\phi}(r)$ and $W_{\phi}(r)$ of $\phi$ mesons are plotted as a function of distance $r$ for the four nuclei ($A=12,16,40$ and $208$). As we observed very small negative mass shift for the $\phi$ mesons, the bound states are not possible for these mesons within the present calculations where the in-medium masses of $K$ and $\bar{K}$ mesons are evaluated using the chiral SU(3) model. In the chiral SU(3) model, the properties of kaons and antikaons are modified differently in the nuclear medium. The effective mass of kaons increases whereas the mass of antikaons decrease as the density of the medium is changed from zero to finite value [65, 66, 64]. In Ref. [6], the $\phi$ meson bound states are investigated considering the modification of kaons and antikaons within the QMC model. The in-medium masses of kaons and antikaons are considered same in the nuclear medium and are observed to decrease as a function of baryon density $\rho_{B}$. Thus, the mass drop in the $\phi$ mesons is also observed to be large as compared to our present calculations. For example, at $\rho_{B}=\rho_{0}$, $\Delta m_{\phi}$ is observed to be $\approx 20$ MeV for the cuttoff $\Lambda_{c}=2000$ MeV. This larger mass drop in the $\phi$ meson masses also leads to the possibility of bound states in the calculations of Ref. [6].
V Summary
In summary, we investigated the possibility of the formation of bound states for neutral mesons $\eta,\eta^{{}^{\prime}}$, $D^{0},\bar{D^{0}}$, $B^{0},\bar{B^{0}}$, $\bar{K^{0}}$ and $\phi$ mesons with nuclei, 12C, 16O, 40Ca and 208Pb. The binding energy for different mesons were calculated by solving the Klein Gordon equation considering the complex potential. The imaginary potential is introduced to understand the impact of absorption of mesons in the nuclei on the formation of bound states. The mass shifts for different mesons considered in the present work are used to obtain the real and imaginary optical potentials for nuclei of different mass numbers. Except for vector $\phi$ mesons, the imaginary optical potential for all other mesons is introduced in the current work through phenomenological prescription in terms of a parameter $\kappa$ following some previous studies [14].
For the study of bound states for pseudoscalar $\eta$ and $\eta^{{}^{\prime}}$ mesons within the chiral SU(3) model, we considered the mixing between the states $\eta_{8}$ and $\eta_{0}$. We observed appreciable mass drop for the $\eta$ mesons whereas $\eta^{{}^{\prime}}$ mass is observed to decrease very less. This leads to the possibility of bound states of $\eta$ mesons with the nuclei considered in this work, whereas for the $\eta^{{}^{\prime}}$ meson no bound state is observed. Using the chiral SU(3) model, generalized to SU(4) sector, we also elaborated on the formation of bound states for $D^{0}$ and $\bar{D^{0}}$ mesons. For both $D^{0}$ and $\bar{D^{0}}$ mesons the mass drop observed at nuclear saturation density is such that it lead to formation of bound states. Among $D^{0}$ and $\bar{D^{0}}$, former is observed to undergo larger mass drop and this lead to the formation of bound states with higher states also. For the case of $B^{0}$ and $\bar{B^{0}}$ mesons also the obtained negative mass shift lead to the formation of bound states for lower and excited states, especially for nuclei with larger mass number $A$. In our present work we also observed that the pseudoscalar $\bar{K^{0}}$ mesons can form the bound states as negative mass shift is observed within the chiral SU(3) model. For the case of vector $\phi$ mesons the mass drop obtained in current work is not large enough to form their bound states with the light and heavy nuclei. The present investigations on the formation of meson-nuclei bound states will be significant for the experiments of $\bar{\text{P}}$ at FAIR, WASA at COSY and J-PARC facilities.